Calculate the volume occupied by 5g of acetylene gas at 50 Celsius and 740 mm pressure

Grade:11

1 Answers

Arun
6 years ago
Dear Siddhartha
 
Applying the general gas equation
 
PV = nRT = m/M RT
 
Here, Mol. wt. of acetylene i.e., C2H2 (M) = 26, P = 740/760 atm,
 
T = 50oC = 50 + 273 = 323 K
 
∴ V = mRT/MP or V 5 * 0.082 *323 *760/26 *740 = 5.23 l
 
Regards
Arun (AskIITians forum expert)